OpenStudy (anonymous):

How would I go about finding the intersections of these two equations? (The first one is a circle and the second is a parabola) (posting equations in comments)

OpenStudy (anonymous):

\[x ^{2}+y ^{2}-16y+39=0\]\[y ^{2}-x ^{2}-9=0\]

OpenStudy (anonymous):

make x the subject of the 2nd equation and insert into the 1st

OpenStudy (anonymous):

so x^2 = y^2 - 9 ?

OpenStudy (anonymous):

yh, sub into 1st

OpenStudy (mathmale):

Tammi: Please hold it. Check your second equation. I think that's the equation of a hyperbola, not a parabola.

OpenStudy (anonymous):

Oh yes, it is! My bad

OpenStudy (anonymous):

And how do I sub into 1st...

OpenStudy (mathmale):

Tammi: Check that 2nd again, please. Solve for y^2. Doesn't matter all that much which oone you solve for, x^2 or y^2, but please solve for y^2 here.

OpenStudy (anonymous):

x^2 +y^2 -16y +39=0 and x^2 = y^2 -9 so (y^2-9) + y^2 - 16y + 39 = 0

OpenStudy (anonymous):

so then -y^2 = -x^2 - 9 ? I don't know how to get rid of the exponent

OpenStudy (mathmale):

Take a moment and ask y ourself why you're solving the 2nd equation for y^2 and why you're substituting the result into the first equation.

OpenStudy (anonymous):

Because there are two ys in that equation and one x ?

OpenStudy (mathmale):

The reason is that substitution eliminates one variable. From the second equation you could get either -x^2=9-y^2 or y^2=9+x^2. Look carefully at the 1st equation presented in this problem and ask yourself which you

OpenStudy (mathmale):

should substitute, y^2=9+x^2 or x^2+y^2-9. Sorry this took so long.

OpenStudy (anonymous):

I'd substitute for y^2 because x^2 is not negative...? So the equation would become \[x ^{2}+(9+x ^{2})-16y+39=0\]?

OpenStudy (mathmale):

Thanks for using the equation editor. As before, my point is that substitution can enable you to eliminate one variable. In the equation y ou've just typed, you still have both x and y in the equation.

OpenStudy (mathmale):

So my suggestion is to take the second equation and transform it into x^2=y^2-9. If you substitute that into the first equation, x will disappear completely and you'll be left with only y in your equation. How much sense does this make?

OpenStudy (anonymous):

That makes perfect sense, so\[(y ^{2}-9)+y ^{2}-16y+39=0\]

OpenStudy (mathmale):

That's really great; I see you know your stuff. Now would y ou please combine those two y^2 terms and type out the equation as it now remains?

OpenStudy (anonymous):

\[2y ^{2}-16y+30=0\]

OpenStudy (mathmale):

Cool, great work. Divide all 3 terms by 2 to siimplify the equation, then type out the new form of the equation.

OpenStudy (anonymous):

\[y ^{2}-8y+15=0\]

OpenStudy (mathmale):

Draw on your past experience to plan what you need to do next to solve for y.

OpenStudy (anonymous):

factor it?

OpenStudy (mathmale):

That might work, and if it does, your suggestion is better than what I had had in mind. Would you try factoring Y^2-8y+15=0.

OpenStudy (anonymous):

(y - 3)(y - 5) (y * y)(y * -5)(-3 * y)(-3 * -5)\[y ^{2}-5y-3y+15\]\[y ^{2}-8y+15\]So (y - 3)(y - 5) are factors

OpenStudy (mathmale):

Great, and so what are the two solutions (y-values)?

OpenStudy (anonymous):

y = 3 and y = 5

OpenStudy (mathmale):

Yes! Now, if you'd please go back to the 2nd equation presented in this problem, and substitute y=3, you'll find that x=0

OpenStudy (mathmale):

What happens if you substitute y=5? You'll get a different answer.

OpenStudy (mathmale):

So, Tammi: if y=3, x=0 and y our solution therre is (0,3). And if y=5, x = ??? and so you'll have 2 more solutions: (???, 5) and (???, 5). Does this make sense to you? As much as I've enjoyed

OpenStudy (mathmale):

working with you, I need to leave. You could check y our results by substituion into the 1st or the origianl equations.

OpenStudy (anonymous):

\[(5)^{2}-x ^{2}-9=0\]\[25 - x ^{2}-9=0\]\[x ^{2} + 16 = 0\]\[x ^{2}=-16\]\[x = \pm4\]?

OpenStudy (mathmale):

Actually, Tammi, you'd get -x^2+16=0. Thus, x would be plus or minus 4, as you have it.

OpenStudy (anonymous):

Oh okay c:

OpenStudy (mathmale):

So, are you OK, are y ou satisfied, or what? Strongly suggest your check your answers by subsitution of the 3 points back into the original (top) equation.

OpenStudy (anonymous):

Yes, I'm in the process of doing that c: I'll post my final answers when I'm done, but I'm okay now (: thank you

OpenStudy (mathmale):

Best of luck to you, Tammi. You know a lot already, and I respect you for it.